OpenStudy (howard-wolowitz):

Can someone help me here:

OpenStudy (howard-wolowitz):

@jdoe0001

OpenStudy (jdoe0001):

\(\textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Upwards \\ \quad \\ {\color{green}{ D}} < 0 & Downwards\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ % template start f(t) = {\color{purple}{ 6}} sin\left( {\color{blue}{ 3}}t - {\color{red}{ \frac{\pi }{6}}} \right) - {\color{green}{ 1}}\\ % template ends \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to\ the\ left \\ \quad \\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to\ the\ right\end{array} \end{array} \\ \quad \\ period\to \cfrac{2\pi }{{\color{blue}{ B}}}\)

OpenStudy (howard-wolowitz):

D

OpenStudy (jdoe0001):

yeap